# Optimal Non-Adaptive Group Testing with One-Sided Error Guarantees

### v3

@@ IV Conclusion
  • $\Phi\coloneqq\log\left(\frac{g^{*}L}{w_{-}}\right)$;
  • $C(L,w_{-})\coloneqq\exp\left(\frac{L^{2}}{4w_{-}}\right)$.
We now proceed to establish (119). We use the following bound derived in [19, Eqn. 37] as a starting point:
$\Lambda$ $\coloneqq\sum_{i\in S}\mathbb{P}(L_{i}=0)$
$\leq\frac{kC(L,w_{-})}{2^{L}}\exp\left((\delta+e^{\Phi}(1-\delta)L)\right)+o(1).$ (120)
Using the definitions of $g^{*}$ and $w_{-}$ along with $L=\frac{T\log 2}{k}$, we can express $\Phi$ as
@@
$\Lambda-o(1)$ $\leq\frac{kC(L,w_{-})}{2^{L}}\exp\left(\Big(\delta+2\log 2\cdot\Big(\frac{n}{k}\Big)^{-\xi}\Big)L\right)$ (126)
$\leq kC(L,w_{-})\exp\Big(-L\log 2+\Big(\delta+2\log 2\cdot\Big(\frac{n}{k}\Big)^{-\xi}\Big)L\Big)$ (127)
$=kC(L,w_{-})\exp\left(-\frac{T\log^{2}2}{k}+\big(\delta+2\log 2\cdot\Big(\frac{n}{k}\Big)^{-\xi}\Big)\frac{T\log 2}{k}\right)$ (128)
@@
$\Lambda$ $\leq 2k\left(\frac{k}{n}\right)^{(1+\varepsilon)(1-\frac{2\delta}{\log 2})}+o(1)$ (132)
$=o(\eta^{-}k),$ (133)
where (133) holds by the assumption that $\eta^{-}=\Omega(k^{-\lambda(\frac{1}{\theta}-1)})$ for some $\lambda\in[0,1)$, provided that $\delta$ is sufficiently small. Denoting the number of items $i\in S$ such that $L_{i}=0$ by $D$, it holds that $\mathbb{E}D=\Lambda$. Using the aforementioned bound on $\Lambda$ and Markov’s inequality, we thus have the following high probability bound on there being $\eta^{-}k$ or more such items:
$P_{\mathrm{e}}^{-}$ $=\mathbb{P}(D\geq\eta^{-}k)=o(1).$ (134)
Additionally, $T=\big(\frac{1}{\log^{2}2}k\log\frac{n}{k}\big)(1+\varepsilon)$ tests being sufficient for this to be the case implies a rate of $\log 2$ is achievable for SUBSET, as desired. ∎
@@
In this appendix, we consider the problem of recovering an estimate $\hat{S}$ that contains at most $\alpha_{1}k$ false negatives and $\alpha_{2}k$ false positives for two fixed and possibly distinct parameters $\alpha_{1},\alpha_{2}\in(0,1)$. While we are not aware of any works giving explicit bounds on the number of tests necessary and sufficient to solve such a problem, we will show that both the achievability and converse bounds follow almost immediately from existing works, with the converse bound being a consequence of [11, Theorem 1] and the achievability bound following from [12, Theorem 1].
@@
[11, Theorem 1] Fix $\alpha_{1}\in(0,1)$ and suppose that the decoder’s list size is $L^{*}\geq k$. Then in order to have $P_{\mathrm{e}}(L^{*},\alpha_{1})\coloneqq\mathbb{P}(\lvert{\mathcal{L}}\cap S\rvert<(1-\alpha_{1}k))\not\to 1$ as $n\to\infty$, we require that
@@
We claim that this converse implies a converse for the asymmetric approximate recovery problem. To see this, we show that achievability in the asymmetric approximate recovery problem implies achievability in the “list approximate recovery” problem (as specified in the previous paragraph), which means that an impossibility result in the list approximate recovery setting is also valid for asymmetric recovery by taking the contrapositive.
Consider an estimate $\hat{S}$ such that $\hat{S}$ has at most $\alpha_{1}k$ false positives and $\alpha_{2}k$ false positives. This then implies that $(1-\alpha_{1})k\leq\lvert\hat{S}\rvert\leq(1+\alpha_{2})k$, since one of the bounds on the number of false positives or negatives would trivially be violated if $\lvert\hat{S}\rvert$ were outside this range. We can then extend this estimate $\hat{S}$ to a list $\mathcal{L}$} such that $\lvert{\mathcal{L}}\rvert=(1+\alpha_{2})k$ by adding in arbitrary items. Since there are at most $\alpha_{1}k$ false negatives in $\hat{S}$ by assumption, the same holds for $\mathcal{L}$} since adding items can only decrease the number of false negatives, implying that $\lvert{\mathcal{L}}\cap S\rvert\geq(1-\alpha_{1})k$. Hence, achievability in the asymmetric approximate recovery setting implies achievability in the list approximate recovery setting. By Theorem 14, if $T\leq(1-\alpha_{1})\big(k\log_{2}\frac{n}{k}\big)(1-\varepsilon)$ for any fixed, arbitrarily small $\varepsilon>0$, then $P_{\mathrm{e}}(L^{*},\alpha_{1})\to 1$. Hence, if $T\leq(1-\alpha_{1})\big(k\log_{2}\frac{n}{k}\big)(1-\varepsilon)$, the error probability for asymmetric approximate recovery also tends to $1$.
@@ Acknowledgment
This work is supported by the National University of Singapore under the Presidential Young Professorship grant scheme. The authors would also like to thank Sidharth Jaggi for helpful discussions surrounding the asymmetric approximate recovery problem.

### v4

@@ IV Conclusion
  • $\Phi\coloneqq\log(\frac{g^{*}L}{w_{-}})$;
  • $C(L,w_{-})\coloneqq\exp(\frac{L^{2}}{4w_{-}})$.
We now proceed to establish (144). We use the following bound derived in [23, Eqn. 37] as a starting point:
$\Lambda$ $\coloneqq\sum_{i\in S}\mathbb{P}(L_{i}=0)$ (145)
$\leq\frac{kC(L,w_{-})}{2^{L}}(1+\delta+e^{\Phi}(1-\delta))^{L}+o(\eta k)$ (146)
$\leq\frac{kC(L,w_{-})}{2^{L}}\exp\left((\delta+e^{\Phi}(1-\delta))L\right)+o(\eta k),$ (147)
where in (146) the $o(\eta k)$ term consists of terms that are exponentially decaying [^9] The terms in question are the last two terms in [23, Eqn. 30]. We refer the interested reader to [23, Eqns. 40-42] and the set of equations above them for explicit bounds on these terms that establish their exponential decay. in $n$ and $k$ (and therefore $o(\eta k)$ for the range of $\eta$ we consider), and (147) uses $1+x\leq e^{x}$. Using the definitions of $g^{*}$ and $w_{-}$ along with $L=\frac{T\log 2}{k}$, we can express $\Phi$ as
@@
$\Lambda-o(\eta k)$
$\leq\frac{kC(L,w_{-})}{2^{L}}\exp\Big(\Big(\delta+2\log 2\cdot\Big(\frac{n}{k}\Big)^{-\xi}\Big)L\Big)$ (153)
$=kC(L,w_{-})\exp\Bigg(-L\log 2$
$\qquad\qquad\qquad\quad+\Big(\delta+2\log 2\cdot\Big(\frac{n}{k}\Big)^{-\xi}\Big)L\Bigg)$ (154)
$=kC(L,w_{-})\exp\Bigg(-\frac{T\log^{2}2}{k}$
$\qquad\quad\quad+\Big(\delta+2\log 2\cdot\Big(\frac{n}{k}\Big)^{-\xi}\Big)\frac{T\log 2}{k}\Bigg)$ (155)
@@
$\Lambda$ $\leq 2k\left(\frac{k}{n}\right)^{(1+\varepsilon)(1-\frac{2\delta}{\log 2})}+o(\eta k)$ (159)
$=o(\eta k),$ (160)
where (160) holds by the assumption that $\eta=\Omega((\frac{k}{n})^{\lambda})$ for some $\lambda\in[0,1)$, provided that $\delta$ is sufficiently small. Denoting the number of items $i\in S$ such that $L_{i}=0$ by $D$, it holds that $\mathbb{E}D=\Lambda$. Using the aforementioned bound on $\Lambda$ and Markov’s inequality, we thus have the following high probability bound on there being $\eta k$ or more such items:
$P_{\mathrm{e}}^{-}$ $=\mathbb{P}(D\geq\eta k)=o(1).$ (161)
Additionally, $T=\big(\frac{1}{\log^{2}2}k\log\frac{n}{k}\big)(1+\varepsilon)$ tests being sufficient for this to be the case implies that a rate of $\log 2$ is achievable for SUBSET, as desired. ∎
@@
In this appendix, we consider the problem of recovering an estimate $\hat{S}$ that contains at most $\alpha_{1}k$ false negatives and $\alpha_{2}k$ false positives for two fixed and possibly distinct parameters $\alpha_{1},\alpha_{2}\in(0,1)$. While we are not aware of any works giving explicit bounds on the number of tests necessary and sufficient to solve such a problem, we will show that both the achievability and converse bounds follow almost immediately from existing works, with the converse bound being a consequence of [29, Theorem 1] and the achievability bound following from [32, Theorem 1].
@@
[29, Theorem 1] Fix $\alpha_{1}\in(0,1)$ and suppose that the decoder’s list size is $L^{*}\geq k$. Then in order to have $P_{\mathrm{e}}(L^{*},\alpha_{1})\coloneqq\mathbb{P}(\lvert{\mathcal{L}}\cap S\rvert<(1-\alpha_{1})k)\not\to 1$ as $n\to\infty$, we require that
@@
We claim that this converse implies a converse for the asymmetric approximate recovery problem. To see this, we show that achievability in the asymmetric approximate recovery problem implies achievability in the “list approximate recovery” problem (as specified in the previous paragraph), which means that an impossibility result in the list approximate recovery setting is also valid for asymmetric recovery (by taking the contrapositive).
Consider an estimate $\hat{S}$ such that $\hat{S}$ has at most $\alpha_{1}k$ false negatives and $\alpha_{2}k$ false positives. This implies that $(1-\alpha_{1})k\leq\lvert\hat{S}\rvert\leq(1+\alpha_{2})k$, since one of the bounds on the number of false positives or negatives would trivially be violated if $\lvert\hat{S}\rvert$ were outside this range. We can then extend this estimate $\hat{S}$ to a list $\mathcal{L}$} such that $\lvert{\mathcal{L}}\rvert=(1+\alpha_{2})k$ by adding in arbitrary items. Since there are at most $\alpha_{1}k$ false negatives in $\hat{S}$ by assumption, the same holds for $\mathcal{L}$} since adding items can only decrease the number of false negatives. Hence, $\lvert{\mathcal{L}}\cap S\rvert\geq(1-\alpha_{1})k$, meaning that achievability in the asymmetric approximate recovery setting implies achievability in the list approximate recovery setting. By Theorem 15, if $T\leq(1-\alpha_{1})\big(k\log_{2}\frac{n}{k}\big)(1-\varepsilon)$ for any fixed, arbitrarily small $\varepsilon>0$, then $P_{\mathrm{e}}(L^{*},\alpha_{1})\to 1$. Hence, if $T\leq(1-\alpha_{1})\big(k\log_{2}\frac{n}{k}\big)(1-\varepsilon)$, the error probability for asymmetric approximate recovery also tends to $1$.
@@ Acknowledgment
This work is supported by the National University of Singapore under the Presidential Young Professorship grant scheme. The authors would like to thank Sidharth Jaggi for helpful discussions surrounding the asymmetric approximate recovery problem.